OpenStudy (anonymous):

find the area of the surface. The part of the sphere x^2+y^2+z^2=64 that lies above the plane z=3

ganeshie8 (ganeshie8):

find the partials and plugin

ganeshie8 (ganeshie8):

\(x^2+y^2+z^2=64\) \(f_x = \dfrac{\partial z}{ \partial x} = \dfrac{-x}{z}\) \(f_y = \dfrac{\partial z}{ \partial y} = \dfrac{-y}{z}\)

ganeshie8 (ganeshie8):

surface area : \[\iint \limits_R \sqrt{f_x^2 + f_y^2 + 1} ~dA\]

ganeshie8 (ganeshie8):

switching to polar may help..

ganeshie8 (ganeshie8):

\[\iint \limits_R \sqrt{\frac{x^2}{z^2} + \frac{y^2}{z^2} + 1} ~dA\] \[\iint \limits_R \sqrt{\frac{x^2+y^2+z^2}{z^2} } ~dA\] \[\iint \limits_R \sqrt{\frac{64}{z^2} } ~dA\] \[8\iint \limits_R \dfrac{1}{\sqrt{64-x^2-y^2}} ~dA\]

ganeshie8 (ganeshie8):

fine, so far ?

OpenStudy (anonymous):

So far I am understanding. Thanks

ganeshie8 (ganeshie8):

next setup the bounds

ganeshie8 (ganeshie8):

the region of interest is above z = 3 and below the sphere : 64-x^2-y^2<= 3^2 x^2 + y^2 <= 55

ganeshie8 (ganeshie8):

so the shadow in xy plane is just a disk of radius sqrt(55), maybe sketch it roughly if u dont see it yet..

OpenStudy (anonymous):

If you do not mind can you guide me all the way to the end of the problem. If you are not to busy please?

ganeshie8 (ganeshie8):

sure :) did u get why the shadow in xy plane is a disk of radius sqrt(55) ?

OpenStudy (anonymous):

yes and thanks

ganeshie8 (ganeshie8):